OpenStudy (anonymous):

There are 9 candidates running for 3 seats on a committee. How many different election results are possible?

OpenStudy (anonymous):

27

OpenStudy (raden):

it means 9C3 = 9! / (9-3)!3! = 9! / 6!3! = 9*8*7*6! / 6!*3*2*1 = 84